# Primal-Dual iLQR

### 13 Mar 2024 v3

@@ I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal component of the Newton-KKT system, and use a dual LQR backward pass to solve its dual component. We also present a new parallel algorithm for solving the dual component of the Newton-KKT system in $O(\log(N))$ parallel time, where $N$ is the number of stages. Combining it with [17], we are able to solve the full Newton-KKT system in $O(\log(N))$ parallel time. The remaining parts of our method have constant parallel time complexity per iteration. Therefore, this paper provides, for the first time, a practical, highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems. As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections or dimension expansions, improving on existing direct multiple shooting approaches such as acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6].
@@ II-B Related Work
Naïf applications of generic optimization methods to unconstrained discrete-time optimal control problems would require $O(N^{3}(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem. The first algorithm to improve on this was DDP [7], achieving complexity $O(N(n+m)^{3})$. Its convergence properties were studied in (Shoemaker and Liao 1991). iLQR [11] can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation. Stagewise Newton [1] exploits the fact that the computation of a Newton step on the “eliminated problem” (obtained by eliminating the variables $x_{i}$ in decreasing order of $i$ by plugging in the dynamics into the costs) can be reduced to solving an LQR problem, thereby achieving the same computational complexity as DDP.
@@
[3] introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above. There are no available convergence results for (iLQR-)GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence. [6] introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes. There are also no available convergence results for FDDP. In both cases, as second order derivatives of the dynamics are not considered, there cannot be local quadratic convergence.
@@ II-C Contributions
We also introduce a new method for parallelizing the computation of the dual solution of the Newton-KKT systems, with parallel time complexity $O(\log(N)\log(n))$, by framing the dual LQR backward pass as an associative scan [18]. This can be combined with the method of [17] to solve the primal-dual Newton-KKT system in parallel time complexity logarithmic in $N$.
@@ IV Algorithm Derivation
In this case, the primal part of the SQP problem is an LQR problem, which can be efficiently solved.
@@
In order to ensure that the matrices $Q_{i}$ are positive semi-definite, it may be helpful to maintain regularization parameters $\mu_{i}$ and always use $Q_{i}+\mu_{i}I$ instead of $Q_{i}$. If, during the LQR process, we detect that $Q_{i}+\mu_{i}I$ is not positive definite, $\mu_{i}$ can be updated by multiplying it by an updated factor $r>1$. When increasing $\mu_{i}$ is not required, we can instead update $\mu_{i}$ by dividing it by $r$. Minimum and maximum regularization parameters may also be established, to allow, when possible, $\mu_{i}$ to eventually be set to 00 if $Q_{i}$ is consistently positive definite, as well as to prevent exploring unreasonably high regularization parameters.
Next, we show how to efficiently compute the $\Delta\lambda$, i.e. how to solve the dual part of the SQP problem. Having determined $\Delta x$, this can be done by specializing $\nabla_{xx}\mathcal{L}(x,\lambda)\Delta x+J(c)(x)^{T}\Delta\lambda=-\nabla_{x}\mathcal{L}(x,\lambda)$. Note that we discard the rows where $J(c)(x)^{T}$ contains any of the $B_{i}$ (as they are not useful for computing the $\Delta\lambda_{i}$), and keep the rows where it contains any of the $A_{i}$.
Letting $l_{i}^{x},l_{i}^{u},Q_{i}^{xx},Q_{i}^{xu},Q_{i}^{ux},Q_{i}^{uu}$ be such that
$l_{i}=\begin{pmatrix}l_{i}^{x}\\ l_{i}^{u}\end{pmatrix},Q_{i}=\begin{pmatrix}Q_{i}^{xx}&Q_{i}^{xu}\\ Q_{i}^{ux}&Q_{i}^{uu}\end{pmatrix},\forall i\in\{0,\ldots,N-1\},$ (46)
it holds that
$\begin{cases}Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}-\Delta\lambda_{i}+A_{i}^{T}\Delta\lambda_{i+1}=\\ -l_{i}^{x},\forall i\in\{0,\ldots,N-1\}\\ Q_{N}\Delta x_{N}-\Delta\lambda_{N}=-l_{N}\end{cases}\Rightarrow$ (47)
$\begin{cases}\Delta\lambda_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+A_{i}^{T}\Delta\lambda_{i+1}+l_{i}^{x},\\ \forall i\in\{0,\ldots,N-1\}\\ \Delta\lambda_{N}=Q_{N}\Delta x_{N}+l_{N}.\end{cases}$
Note that these equations can be used to recursively compute the $\Delta\lambda_{i}$, in decreasing order of $i$. We call this method the dual LQR backward pass.
Moreover, note that these $\Delta\lambda_{i}$ also satisfy the block-rows of $\nabla_{xx}\mathcal{L}(x,\lambda)\Delta x+J(c)(x)^{T}\Delta\lambda=-\nabla_{x}\mathcal{L}(x,\lambda)$ that we discarded, as a solution to the full Newton-KKT system must exist due to the LICQ conditions being satisfied, and as these $\Delta\lambda_{i}$ are the only solution that satisfies the block-rows we did not discard.
Using the methods described above, we can efficiently compute both $\Delta x$ and $\Delta\lambda$. The remaining parts of the optimization algorithm we describe here don’t require further specialization in the case of unconstrained discrete-time optimal control problems.
V Parallel Dual LQR Backward Pass with Associative Scans
@@
In this section, we give a brief description of what associative scans are, and how they can also be used to derive a similar method that is equivalent to the dual LQR backward pass, while only taking $O(\log(N)\log(n))$ parallel time. Note that the $\log(n)$ factor here comes from the parallel time complexity of taking the product of two matrices of dimension $n$.
@@
As shown in eq. 47, the dual solutions $\Delta\lambda_{i}$ of the Newton-KKT system eq. 10 can be obtained by applying a backward pass starting at $\Delta\lambda_{N}=Q_{N}\Delta x_{N}+l_{N}$, with recursion rule $\Delta\lambda_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+A_{i}^{T}\Delta\lambda_{i+1}+l_{i}^{x}$. We can simplify the notation by defining, for $i\in\{0,\ldots,N-1\}$, $M_{i}=A_{i}^{T}$ and $m_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+l_{i}^{x}$, so that the recursion rule becomes simply $\Delta\lambda_{i}=M_{i}\Delta\lambda_{i+1}+m_{i}$.
Moreover, we define $m_{N}=Q_{N}\Delta x_{N}+l_{N}$ and $M_{N}=I$.
Letting $\mathcal{X}=\mathbb{R}^{n}\times\mathbb{R}^{n\times n}$ and letting $f:\mathcal{X}\rightarrow\mathcal{X}$ be defined by $f((a,B),(c,D))=(a+Bc,BD)$, we claim that $f$ is associative and that $\Delta\lambda_{0},\ldots,\Delta\lambda_{N}$ can be obtained by doing a reverse associative scan with respect to $f$ on $(m_{0},M_{0}),\ldots,(m_{N-1},M_{N-1}),(m_{N},M_{N})$.
The associative property is simple to verify:
@@
Let $F$ denote the operator that extracts the first element from its input sequence. We can show by induction that
$(\Delta\lambda_{i},M_{i}\cdots M_{N})=F(S_{r}((m_{i},M_{i})),\ldots,(m_{N},M_{N});f).$ (49)
This is clearly true for $i=N$, as
$S_{r}((m_{N},M_{N});f)=(m_{N},M_{N})=(\Delta\lambda_{N},M_{N}).$ (50)
For $i<N$, by definition of $S_{r}$ and the induction hypothesis,
$F(S_{r}((m_{i},M_{i}),\ldots,(m_{N},M_{N});f))$ (51)
$=$ $f((m_{i},M_{i}),F(S_{r}((m_{i+1},M_{i+1}),\ldots,(m_{N},M_{N});f)))$
$=$ $f((m_{i},M_{i}),(\Delta\lambda_{i+1},M_{i+1}\cdots M_{N}))$
$=$ $(m_{i}+M_{i}\Delta\lambda_{i+1},M_{i}\cdots M_{N})$
$=$ $(\Delta\lambda_{i},M_{i}\cdots M_{N}).$
VI Algorithm Overview

### 09 Apr 2024 v4

@@ I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal-dual Newton-KKT system. As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections or dimension expansions, improving on existing direct multiple shooting approaches such as acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6]. As our algorithm avoids sequential rollouts of the nonlinear dynamics, it can be combined with [17] to run in $O(\log(N))$ parallel time per iteration (where $N$ is the number of stages), as well as $O(1)$ parallel time per line search iteration. Therefore, this paper provides a practical, theoretically sound, and highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems.
@@ II-B Related Work
Naïf applications of generic optimization methods to unconstrained discrete-time optimal control problems would require $O(N^{3}(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem. The first algorithm to improve on this was DDP [7], achieving complexity $O(N(n+m)^{3})$. Its convergence properties were studied in [10]. iLQR [11] can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation. Stagewise Newton [1] exploits the fact that the computation of a Newton step on the “eliminated problem” (obtained by eliminating the variables $x_{i}$ in decreasing order of $i$ by plugging in the dynamics into the costs) can be reduced to solving an LQR problem, thereby achieving the same computational complexity as DDP.
@@
[3] introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above. There are no available convergence results for iLQR-GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence. [6] introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes. There are also no available convergence results for FDDP. In both cases, as second order derivatives of the dynamics are not considered, there cannot be local quadratic convergence.
@@ II-C Contributions
Moreover, we show that the primal-dual Newton-KKT systems we pose can be solved in $O(\log(N)\log(n))$ parallel time complexity, using a minor variation of [18], and that the remaining parts of our method have constant parallel time complexity.
@@ IV Algorithm Derivation
In this case, the the SQP problem is a primal-dual LQR problem, which can be efficiently solved.
@@
In order to ensure that the matrices $Q_{i}$ are positive semi-definite, it may be helpful to maintain regularization parameters $\mu_{i}$ and always use $Q_{i}+\mu_{i}I$ instead of $Q_{i}$. If, during the LQR process, we detect that $Q_{i}+\mu_{i}I$ is not positive definite, $\mu_{i}$ can be updated by multiplying it by an updated factor $r>1$. When increasing $\mu_{i}$ is not required, we can instead update $\mu_{i}$ by dividing it by $r$. Minimum and maximum regularization parameters may also be established, to allow, when possible, $\mu_{i}$ to eventually be set to 00 if $Q_{i}$ is consistently positive definite, as well as to prevent exploring unreasonably high regularization parameters. Another option would be to regularize the $Q_{i}$ by performing an explicit eigenvalue decomposition and removing negative (or non-positive, when positive definiteness is required) eigenvalues.
V Sequential Primal LQR Overview
In this section, we will go over the conventional sequential algorithm for solving primal LQR problems. In the interest of not deviating from standard notation, variable names may henceforth not match earlier parts of this paper. A linear-quadratic regulator (LQR) problem is an unconstrained discrete-time optimal control problem where the costs are quadratic functions and the dynamics are affine functions. Specifically, they are optimization problems of the form
$\min_{x_{0},u_{0}\ldots,x_{N}}$ $\sum\limits_{i=0}^{N-1}\left(\frac{1}{2}x_{i}^{T}Q_{i}x_{i}+q_{i}^{T}x_{i}+\frac{1}{2}u_{i}^{T}R_{i}u_{i}+r_{i}^{T}u_{i}\right.$ (42)
$+x_{i}^{T}M_{i}u_{i}\right)+\frac{1}{2}x_{N}^{T}Q_{N}x_{N}+q_{N}^{T}x_{N}$
s.t. $x_{0}=s_{0}$
$x_{i+1}=A_{i}x_{i}+B_{i}u_{i}+c_{i},\forall i\in\{0,\ldots,N-1\}.$
Note that the $R_{i}$ are required to be positive definite, and that the $Q_{i}-M_{i}R_{i}^{-1}M_{i}^{T}$ are required to be positive semi-definite, otherwise a minimum may not exist.
The conventional method for solving LQR problems sequentially relies on the fact that the optimal cost-to-go functions are quadratic on the state at the corresponding stage. Therefore, such functions $V_{i}(x_{i})=0.5x_{i}^{T}P_{i}x_{i}+p_{i}^{T}x_{i}+z_{i}$ can be computed in decreasing order of stage (i.e. $i$). Moreover, the optimal controls to be applied at each stage can be shown to be affine functions of the state at the corresponding stage, i.e. $u_{i}=K_{i}x_{i}+k_{i}$. The $K_{i}$ and $k_{i}$ can be computed as part of the same backward pass as the $P_{i}$ and $p_{i}$. Note that the constants $z_{i}$ do not need to be computed. Initializing $P_{N}=Q_{N}$ and $p_{N}=q_{N}$, and introducing $G_{i},H_{i},h_{i}$ as convenient auxiliary variables, the following recursion rules can be used:
$G_{i}=R_{i}+B_{i}^{T}P_{i+1}B_{i}$ (43)
$H_{i}=B_{i}^{T}+P_{i+1}A_{i}+M_{i}^{T}$
$h_{i}=B_{i}^{T}(p_{i+1}+P_{i+1}c_{i})+r_{i}$
$K_{i}=-G_{i}^{-1}H_{i}$
$k_{i}=-G_{i}^{-1}h_{i}$
$P_{i}=Q_{i}+A_{i}^{T}P_{i+1}A_{i}+K_{i}^{T}H_{i}$
$p_{i}=q_{i}+A_{i}^{T}(p_{i+1}+P_{i+1}c_{i})+K_{i}^{T}h_{i}$
Once the $K_{i},k_{i}$ have been computed, the $u_{i}$ can be computed in increasing order of $i$, by alternating evaluations of $u_{i}=K_{i}x_{i}+k_{i}$ and $x_{i+1}=A_{i}x_{i}+B_{i}u_{i}+c_{i}$. This is usually called the LQR forward pass.
VI Associative Scans Overview
@@
VII Parallel Primal LQR Overview
A parallel algorithm for solving these problems was presented in [17], although we require some minor changes due to a difference in problem formulations. In order to describe this method, we need to introduce an important definition. An interval value function $V_{i\rightarrow j}(x_{i},x_{j})$ maps states $x_{i},x_{j}$ at stages $i,j$ to the minimum possible cost incurred in stages $i,\ldots,j-1$ among all trajectories that start at state $x_{i}$ in stage $i$ and end at state $x_{j}$ in stage $j$ ($\infty$ if no such trajectory exists). The key insight of [17] is that the optimal interval value functions can be represented as
$V_{i\rightarrow j}(x_{i},x_{j})=\max\limits_{\lambda}\left(\frac{1}{2}x_{i}^{T}P_{i\rightarrow j}x_{i}+p_{i\rightarrow j}^{T}x_{i}\right.$ (44)
$\left.-\frac{1}{2}\lambda^{T}C_{i\rightarrow j}\lambda-\lambda^{T}\left(x_{j}-A_{i\rightarrow j}x_{i}-c_{i\rightarrow j}\right)\right).$
Initializing, for $i\in\{0,\ldots,N-1\}$, $V_{i\rightarrow i+1}$ with
$P_{i\rightarrow i+1}$ $=Q_{i}-M_{i}R_{i}^{-1}M_{i}^{T},$ (45)
$p_{i\rightarrow i+1}$ $=q_{i}-M_{i}R_{i}^{-1}r_{i},$
$A_{i\rightarrow i+1}$ $=A_{i}-B_{i}R_{i}^{-1}M_{i}^{T},$
$C_{i\rightarrow i+1}$ $=B_{i}R_{i}^{-1}B_{i}^{T},$
$c_{i\rightarrow i+1}$ $=c_{i}-B_{i}R_{i}^{-1}r_{i},$
and initializing $V_{N\rightarrow N+1}$ with
$P_{N\rightarrow N+1}$ $=Q_{N},$ (46)
$p_{N\rightarrow N+1}$ $=q_{N},$
$A_{N\rightarrow N+1}$ $=0,$
$C_{N\rightarrow N+1}$ $=0,$
$c_{N\rightarrow N+1}$ $=0,$
the following combination rules can be applied to compute $V_{i\rightarrow k}$ from $V_{i\rightarrow j}$ and $V_{j\rightarrow k}$:
$P_{i\rightarrow k}$ $=A_{i\rightarrow j}^{T}\left(I+P_{j\rightarrow k}C_{i\rightarrow j}\right)^{-1}P_{j\rightarrow k}A_{i\rightarrow j}+P_{i\rightarrow j},$ (47)
$p_{i\rightarrow k}$ $=A_{i\rightarrow j}^{T}\left(I+P_{j\rightarrow k}C_{i\rightarrow j}\right)^{-1}\left(p_{j\rightarrow k}+J_{j\rightarrow k}c_{i\rightarrow j}\right)$
$+p_{i\rightarrow j},$
$A_{i\rightarrow k}$ $=A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}A_{i\rightarrow j},$
$C_{i\rightarrow k}$ $=A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}C_{i\rightarrow j}A_{j\rightarrow k}^{T}+C_{j\rightarrow k},$
$c_{i\rightarrow k}$ $=A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}\left(c_{i\rightarrow j}-C_{i\rightarrow j}p_{j\rightarrow k}\right)$
$+c_{j\rightarrow k}.$
A reverse associative scan [18] can be used to compute the $P_{i\rightarrow N+1},p_{i\rightarrow N+1}$ (i.e. the $P_{i},p_{i}$ from section V), which in turn can be used to compute the $K_{i},k_{i}$ in $O(1)$ parallel time.
Finally, we wish to compute the $x_{i},u_{i}$ from the $K_{i},k_{i}$. Note that the sequential LQR forward pass has $O(T)$ parallel time complexity. However, as done in [17], we can reduce the computation of the $x_{i}$ to a sequential composition of affine functions, which can also be parallelized with an associative scan [18]. This will be described in the next section.
$x_{i+1}$ $=A_{i}x_{i}+B_{i}u_{i}+c_{i}=A_{i}x_{i}+B_{i}\left(K_{i}x_{i}+k_{i}\right)+c_{i}$ (48)
$=\left(A_{i}+B_{i}K_{i}\right)x_{i}+\left(B_{i}k_{i}+c_{i}\right)$
Once these affine functions have been composed, they can be independently applied to $x_{0}$ to recover all the $x_{i}$. The $u_{i}$ can then be computed in $O(1)$ parallel time by independently evaluating $u_{i}=K_{i}x_{i}+k_{i}$.
VIII Composing Affine Functions with Associative Scans
In this section, we describe an $O(\log(N))$ parallel time algorithm for composing $N$ affine functions $F_{i}(x)=M_{i}x_{i}+m_{i}$, as done in [17].
Letting $\mathcal{X}=\mathbb{R}^{n}\times\mathbb{R}^{n\times n}$ and letting $f:\mathcal{X}\rightarrow\mathcal{X}$ be defined by $f((a,B),(c,D))=(Da+c,DB)$, we claim that $f$ is associative. This is simple to verify:
@@ VIII Composing Affine Functions with Associative Scans
Moreover, note that $f$ is the affine function composition operator, as $D(Bx+a)+c=(DB)x+(Da+c)$.
IX Dual LQR Backward Pass
Next, we show how to compute the multipliers $\lambda$, i.e. how to solve the dual part of the LQR problem. Note that primal-dual LQR problems can be solved in a single Newton step, independently of the starting values of $x,\lambda$. Therefore, starting with $x,\lambda=0$, the $\Delta x,\Delta\lambda$ from the Newton-KKT system correspond to the optimal $x,\lambda$. Having determined $x$ using one of the primal LQR algorithms, determining $\lambda$ can be done by specializing $\nabla_{xx}\mathcal{L}(0)x+J(c)(0)^{T}\lambda=-\nabla_{x}\mathcal{L}(0)$ (i.e. the first block-row of the Newton-KKT system). Note that we discard the rows where $J(c)(x)^{T}$ contains any of the $B_{i}$ (as they are not useful for computing the $\lambda_{i}$), and keep the rows where it contains any of the $A_{i}$.
Specifically, for $i\in\{0,\ldots,N-1\}$, it holds that
$Q_{i}x_{i}+M_{i}u_{i}-\lambda_{i}+A_{i}^{T}\lambda_{i+1}=-q_{i}\Rightarrow$ (50)
$\lambda_{i}=Q_{i}x_{i}+M_{i}u_{i}+A_{i}^{T}\lambda_{i+1}+q_{i},$
and that
$Q_{N}x_{N}-\lambda_{N}=-q_{N}\Rightarrow$ (51)
$\lambda_{N}=Q_{N}x_{N}+q_{N}.$
Note that these equations can be used to recursively compute the $\lambda_{i}$, in decreasing order of $i$.
Moreover, note that these $\lambda_{i}$ also satisfy the block-rows of $\nabla_{xx}\mathcal{L}(0)x+J(c)(0)^{T}\lambda=-\nabla_{x}\mathcal{L}(0)$ that we discarded, as a solution to the full Newton-KKT system must exist due to the LICQ conditions being satisfied, and as these $\lambda_{i}$ are the only solution that satisfies the block-rows we did not discard.
X Parallel Dual LQR
While the dual LQR backward pass described in section IX can be parallelized with a reverse associative scan (in a similar fashion to section VIII), we can actually compute the multipliers $\lambda$ in $O(1)$ parallel time.
The multiplier $\lambda_{0}$ is also the Lagrange/KKT multiplier of
$\min_{x_{0}}\frac{1}{2}x_{0}^{T}P_{0}x_{0}+p_{0}^{T}x_{0}\quad\mbox{s.t.}\quad x_{0}=s_{0}.$ (52)
By the first order optimality conditions, the gradient of the corresponding Lagrangian
$\mathcal{L}(x_{0},\lambda_{0})=\frac{1}{2}x_{0}^{T}P_{0}x_{0}+p_{0}^{T}x_{0}+\lambda^{T}(s_{0}-x_{0})$ (53)
must be 00. In particular,
$\nabla_{x_{0}}\mathcal{L}(x_{0},\lambda_{0})=0\Rightarrow\lambda_{0}=P_{0}x_{0}+p_{0}.$ (54)
Similarly, each multiplier $\lambda_{i}$ is also the Lagrange/KKT multiplier of
$\min_{u_{i},x_{i+1}}$ $\frac{1}{2}u_{i}^{T}R_{i}u_{i}+r_{i}^{T}u_{i}+x_{i}^{T}M_{i}u_{i}$ (55)
$+\frac{1}{2}x_{i+1}^{T}P_{i+1}x_{i+1}+p_{i+1}^{T}x_{i+1}$
s.t. $x_{i+1}=A_{i}x_{i}+B_{i}u_{i}+c_{i}.$
The corresponding Lagrangian is
$\mathcal{L}(u_{i},x_{i+1},\lambda_{i+1})$ $=\frac{1}{2}u_{i}^{T}R_{i}u_{i}+r_{i}^{T}u_{i}+x_{i}^{T}M_{i}u_{i}$ (56)
$+\frac{1}{2}x_{i+1}^{T}P_{i+1}x_{i+1}+p_{i+1}^{T}x_{i+1}$
$+\lambda_{i+1}^{T}\left(A_{i}x_{i}+B_{i}u_{i}+c_{i}-x_{i+1}\right).$
Similarly, the first order optimality conditions imply
$\nabla_{x_{i+1}}\mathcal{L}(u_{i},x_{i+1},\lambda_{i+1})=0\Rightarrow\lambda_{i+1}=P_{i+1}x_{i+1}+p_{i+1}.$ (57)
Finally, note that all the $\lambda_{i}$ can be evaluated independently, so a backward pass is not required using these formulas. This method of computing the multipliers also has the advantage of involving fewer computations and being more numerically stable.
XI Algorithm Overview